Find an equation in standard from for this line.
A line with a slope of -2 that goes through the point (1,-3)

Answer by stanbon(48558)

Find an equation in standard from for this line.
A line with a slope of -2 that goes through the point (1,-3)
---------
y=mx+b
-3=(-2)(1)+b
b=-1
y=-2x-1
2x+y=1
Cheers,
Stan H.